OpenStudy (anonymous):

Area of Rhombus is 28m^2 and one of its diagonal is 4m.its perimeter is. a.)4root53 b.)36m c.)2root53 d.)none of the above

Directrix (directrix):

a.)4root53

jhonyy9 (jhonyy9):

D1 =4 --- diagonal 1 and D2 is diagonal second A =28 so A=2(D1*D2/2) 28=4*D2/2 28=2D2 D2 =14 lenght of a side =sqrt(49+4)=sqrt53 so the perimeter of this rhombus will be 4*sqrt53 - a. is right if there 4root53 sign 4sqrt53 what is 4 squarroot53

jhonyy9 (jhonyy9):

hope so much that is understandably sure

jhonyy9 (jhonyy9):

good luck bye